OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (dumbcow):

first divide the "-3" over |x+7| = 4 split into 2 cases x+7 = 4 x+7 = -4 then solve for "x" for each equation

OpenStudy (nurali):

-3|x + 7| = -12 Divide by -3 |x + 7| = 4 x+7=4 or x+7=-4 x=4-7 or x=-4-7 x=-3 or x=-11

OpenStudy (anonymous):

ok